OpenStudy (anonymous):

when (67^(67) + 67) is divided by 68 the remainder is 1) 1 2) 63 3) 66 4) 67

OpenStudy (lgbasallote):

calculator?

OpenStudy (anonymous):

Calculator will not work.......... U need super computer or a super brain...lol

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

...or wolframalpha....

OpenStudy (anonymous):

YEP

OpenStudy (lgbasallote):

you can do it manually actually... 7^1 = 7 7^2 = 49 7^3 = xxx3 7^4 = xxxx1 7^5 = xxxx7 so in every four powers, there is a cycle you have 7^67 67/4 = 16 R 3 so 7^67 has a last digit of 3 if you add that to 67...then the last digit is 0 now...when you divide that by 8...i have no idea what the remainder is

OpenStudy (anonymous):

40 is actually exactly divisible by 8

OpenStudy (anonymous):

@mukushla could help...... but is offline

OpenStudy (anonymous):

I think @experimentX can also do this

OpenStudy (anonymous):

any formulas solve this?

OpenStudy (anonymous):

here is the answer :) http://en.allexperts.com/q/Algebra-2061/2010/6/math-88.htm

OpenStudy (experimentx):

67^5 mod 68 = 67 67^(5*13+2) mod 67 = 67^(1+2) mod 68 with this you get (67^3+67) mod 68 = (67^2 + 1)*67 mod 68 67^2 mod 68 = 67*67 mod 68 = 68 - 67 n*67 mod 68 = 68 - n, if n<68 (67^2 + 1) mod 68 = 2 so we have (67^2 + 1)*67 mod 68 = 2*67 mod 68 = 68 - 2 = 66

OpenStudy (anonymous):

\[67^{67}+67\equiv(-1)^{67}+67\equiv-1+67\equiv66 \ \ \text{mod} \ 68\]

OpenStudy (experimentx):

i can't beat you on algebra man!!

OpenStudy (anonymous):

;)